# A Geometric Tension Dynamics Model of Epithelial Convergent Extension

### v2

@@ 4.4 Tissue scale simulation
For the first half of the work, Figs. 2-5, we use a tissue made of identical cells with free boundary conditions. We now specify the boundary conditions, and the division of the tissue into active and passive patches of the simulations for Figs 7-8.
We use appropriate boundary conditions to mimic the geometry of the Drosophila lateral ectoderm. We consider a initially rectangular tissue patch with height $>$ width. The top/bottom edges correspond to the dorsal pole, the middle to the ventral region, and the left/right sides to the anterior/posterior edges of the lateral ectoderm. To simulate the cylindrical geometry of the Drosophila trunk, the top/bottom boundary are given slip-wall boundary conditions, i.e. the cells on the boundary are pinned along the $y$-axis and free to move along the $x$-axis. The left/right boundaries are free. Boundary conditions are numerically implemented by a confining potential which penalizes deviation of the $x$ or $y$ coordinate of the cell centroids designated to be boundary cells from the location of the slip wall. This allows for flexible implementation of various boundaries and geometries. We can also chose different boundary conditions. We can also add slip walls on the left/right boundaries to model the experiments carried out in Ref. [8], where GBE was blocked by cauterization of the ventral tissue close to the posterior pole.
Next, we designate active and passive regions in the tissue. In the active region, tensions evolve in time according to excitable tension feedback Eq. (S25). In the passive regions, tensions are governed by homeostasis (see below), and cells are softer. To simulate a wild type embryo, we chose the size of the passive patch to be approximately 40% of the embryo circumference (the exact value does not have an important influence). To mimic an embryo without modulation of activity along the DV axis (e.g. a TollRM9 mutant), all cells along the embryo were taken to be active. The cells at the very tissue boundary are always passive for reasons of numerical stability. Otherwise edges on the boundary could undergo a T1, which leads to an invalid state.
We can take model a situation where the cells in simulated tissue are initially stretched isogonally. This can be encoded by stretching the tissue using the slip wall boundary conditions; instead, we set the cell reference shape $S_{0}$ to an anisotropic matrix, as discussed above in the SI. For the simulations of the Fig. 7 and 7, we set the initial stretching to 0.
@@
where $\tau_{h}$ is the homeostasis timescale, and $\tilde{T}_{ij,0}$ is the tension on the edge at the beginning of the simulation. For newly generated edges after a passive T1, we generate a random reference value $\tilde{T}_{ij,0}\approx 1$. One can also set $\tilde{T}_{ij,0}\equiv 1$ for all passive cells, in which case the passive tissue relaxes to a hexagonal lattice. The speed of homeostasis $\tau_{h}$ has no strong influence on the simulations, but must be large enough so that tensions relax after a passive T1 process, before a second passive T1 process takes place. In our simulations, we always set $\tau_{h}=\tau_{p}$, the speed of passive tension relaxation in the active tissue.
@@
For each measurement show in the figures of this paper, by default, we pooled or averaged 3 simulation runs with identical parameters, but different random seeds for the creation of the initial tension triangulation. The phase diagram Fig. 3A shows the average of 6 simulation runs each. The trajectories in Fig. 7 as well as the interface length and orientation statistics in Fig. 8 were computed from a single simulation.
Unless specified otherwise, we simulated cell arrays with $N_{\mathrm{cells}}\approx 10^{3}$ cells (the exact number could vary slightly due to randomness in the hard disk point process). Simulations of saturating feedback in Fig. 5 were performed with $N_{\mathrm{cells}}\approx 350$ cells, and simulations in Figs. 7-8 had $N_{\mathrm{cells}}\approx 750$ cells. Simulation edges were always excluded from analysis. For the single-triangle simulation in Fig. 5, we simulated $4\times 10^{3}$ individual triangles from the initial condition of the full numerical simulations. When showing time traces, we smoothed the T1-rate and aspect ratio time series with a Gaussian kernel of width $10dt$ where $dt=\tau_{T}/200$ is the simulation time step. Numerical parameters were adjusted so that elastic energy minimization and triangulation flattening converged for all simulations and all time points.
Simulations were carried out on a Supermicro workstation with 64 Intel Xeon Gold 6326 cores and 256GB of RAM. However, individual simulations (rather than parameter scans and phase diagrams) can be run without problem on a consumer grade laptop.
@@
Passive region shear moduli $\mu_{p},\>\lambda_{p}$ $0.2$ Higher shear resistance of passive tissue lowers net amount of CE.
@@
Table 1: Parameters of tissue scale simulation. A value in the second-to-last column indicates a “default” value which was used in all simulations except otherwise indicated. $\overline{T}$ is the overall tension scale (average edge tension at initial condition), and $\ell_{0}$ the overall length scale (average edge length at initial condition). Net CE denotes the amount of convergent extension as tissue flow saturates.
5 Movie captions
Movie 1: Tissue shape change by dynamics tension constraints.
Implementation of our model for a symmetric, regular cell array, characterized by one angle $\phi$, determined by the tensions $T_{0},T_{1}$, and two lengths, $\ell_{0}$ and $\ell_{1}$, parametrizing the soft isogonal modes. The contour plot shows the cell shape energy $E_{\mathcal{C}}$ in the incompressible limit where $\ell_{1}$ is determined uniquely by $\ell_{0}$ and $\phi$. Relaxation of the sub-dominant cell shape energy $E_{\mathcal{C}}$ is constrained to the isogonal subspace (white line) determined by tension force balance. For the critical tension ratio $T_{0}/T_{1}=\sqrt{2}$, constraining the angle to $\phi_{c}=\pi/2$ (black dashed line), the interface length minimizing $E_{\mathcal{C}}$ vanishes (red half-disk).
Movie 2: Simulations comparing ordered and disordered initial cell packings.
The total extend of tissue shape change depends on the initial degree of order in the cell packing (cf. Fig. 2). Starting from an hexagonal packing, cell rearrangements are initially well-coordinated leading to coherent tissue flow. As order is lost, tissue flow stalls. The number of cells and the initial tension anisotropy, as well as all other parameters, are the same for both simulations. The simulations were carried out with free boundary conditions. Top row shows the cell arrays, colored by cell coordination number (hexagons=white), bottom row the underlying tension triangulation.
Movie 3: Simulations comparing saturating and winner-takes-all tension feedback.
Comparison of saturating (left) and winner-takes-all (right) tension feedback (cf. Fig. 5). Both simulations were initialized identical initial conditions (an ordered cell array with 20% tension anisotropy). All other parameters are the same for both simulations. Top row shows the cell arrays, colored by cell coordination number (hexagons=white), bottom row the underlying tension triangulation. Saturating feedback leads to little large scale shape change and creates tension cables.
Movie 4: Simulation combining active and passive tissue regions.
The embryo is modeled by a rectangular tissue patch with slip walls on the top and bottom edge, mimicking the embryo’s cylindrical geometry (cf. Fig. 7). The active and passive tissue regions, indicated by color and shading of the edges (color/solid=active), correspond to the germ band and the amnioserosa, respectively. Left: Cell tessellation with black dots indicating the cells along the slip wall (restricted to move along horizontally along the “AP axis”). Right: underlying tension triangulation. The tension triangulation is initialized with the experimentally observed initial tension anisotropy which breaks rotational symmetry and aligns the tissue extension along the “AP axis”. Colored dots follow a selected subset cells to highlight rearrangements.
Movie 5: Simulation of convergent extension without cell rigidity and without DV modulation of activity.
Simulation of tissue scale model of convergent extension in an embryo-like geometry for a control and two “mutant” configurations, corresponding Fig. 8. Left: “Control”, as in Movie 3. Middle: Active tissue with vanishing shear modulus $\mu_{a}=0$. Right: No DV modulation, i.e. no soft, passive region. All other parameters are identical. Top shows cell array, bottom underlying tension triangulation. Active and passive tissue regions are indicated by color and shading of the edges (color/solid=active). In all three conditions, active T1s drive elongation of the tension triangulation, while the convergent extension of the cell array only occurs for the “control” condition.

### v3

@@ 5.4 Tissue scale simulation
For the first half of the work, Figs. 24, we use a tissue made of identical cells with free boundary conditions. We now specify the boundary conditions, and the division of the tissue into active and passive patches of the simulations for Figs. 67.
We use appropriate boundary conditions to mimic the geometry of the Drosophila lateral ectoderm. We consider an initially rectangular tissue patch with height $>$ width. The top/bottom edges correspond to the dorsal pole, the middle to the ventral region, and the left/right sides to the anterior/posterior edges of the lateral ectoderm. To simulate the cylindrical geometry of the Drosophila trunk, the top and bottom boundaries are given slip-wall boundary conditions, i.e. the cells on the boundary are pinned along the $y$-axis and free to move along the $x$-axis. The left/right boundaries are free. Boundary conditions are numerically implemented by a confining potential which penalizes deviation of the $x$ or $y$ coordinate of the cell centroids designated to be boundary cells from the location of the slip wall. This allows for flexible implementation of various boundaries and geometries. We can also choose different boundary conditions. We can also add slip walls on the left/right boundaries to model the experiments carried out in Ref. [Collinet.etal2015], where GBE was blocked by cauterization of the ventral tissue close to the posterior pole.
Next, we designate active and passive regions in the tissue. In the active region, tensions evolve in time according to excitable tension feedback (S26). In the passive regions, tensions are governed by homeostasis (see below), and cells are softer. To simulate a wild-type embryo, we chose the size of the passive patch to be approximately 40% of the embryo circumference (the exact value does not have an important influence). To mimic an embryo without modulation of activity along the DV axis (e.g. a TollRM9 mutant), all cells along the embryo were taken to be active. The cells at the very tissue boundary are always passive for reasons of numerical stability. Otherwise, edges on the boundary could undergo a T1, which leads to an invalid state.
We model a situation where the cells are initially stretched isogonally. We implement this scenario by making the reference cell shape $S_{0}$ anisotropic, as discussed above. For the simulations shown in Figs. 6 and 7, we set the initial stretching to zero.
@@
where $\tau_{h}$ is the homeostasis timescale, and $\tilde{T}_{ij,0}$ is the tension on the edge at the beginning of the simulation. For newly generated edges after a passive T1, we generate a random reference value $\tilde{T}_{ij,0}\approx 1$. One can also set $\tilde{T}_{ij,0}\equiv 1$ for all passive cells, in which case the passive tissue relaxes to a hexagonal lattice. The speed of homeostasis $\tau_{h}$ has no strong influence on the simulations, but must be large enough so that tensions relax after a passive T1 process before a second passive T1 process takes place. In our simulations, we always set $\tau_{h}=\tau_{p}$, the speed of passive tension relaxation in the active tissue.
@@
For each quantification shown in the figures of this paper, by default, we pooled or averaged 3 simulation runs with identical parameters while using different random seeds for the creation of the initial tension triangulation. The phase diagram Fig. 2D shows the average of 6 simulation runs each. The time traces in Fig. S10B as well as the interface length and orientation statistics in Fig. S10C were computed from a single simulation.
Unless specified otherwise, we simulated cell arrays with $N_{\mathrm{cells}}\approx 10^{3}$ cells (the exact number could vary slightly due to randomness in the hard disk point process). Simulations of saturating feedback in Fig. 5 were performed with $N_{\mathrm{cells}}\approx 350$ cells, and simulations in Figs. 67 had $N_{\mathrm{cells}}\approx 750$ cells. Simulation edges were always excluded from the analysis. For the single-triangle simulation in Fig. 4, we simulated $4\times 10^{3}$ individual triangles from the initial condition of the full numerical simulations. When showing time traces, we smoothed the T1-rate and aspect ratio time series with a Gaussian kernel of width $10dt$ where $dt=\tau_{T}/200$ is the simulation time step. Numerical parameters were adjusted so that elastic energy minimization and triangulation flattening converged for all simulations and all time points.
Simulations were carried out on a Supermicro workstation with 64 Intel Xeon Gold 6326 cores and 256GB of RAM. However, individual simulations (rather than parameter scans and phase diagrams) can be run without problem on a consumer-grade laptop.
@@
6 Movie Captions
Tissue shape change by dynamics tension constraints. Implementation of our model for a symmetric, regular cell array, characterized by one angle $\phi$, determined by the tensions $T_{0},T_{1}$, and two lengths, $\ell_{0}$ and $\ell_{1}$, parametrizing the soft isogonal modes. The contour plot shows the cell shape energy $E_{\mathcal{C}}$ in the incompressible limit where $\ell_{1}$ is determined uniquely by $\ell_{0}$ and $\phi$. Relaxation of the sub-dominant cell shape energy $E_{\mathcal{C}}$ is constrained to the isogonal subspace (white line) determined by tension force balance. For the critical tension ratio $T_{0}/T_{1}=\sqrt{2}$, constraining the angle to $\phi_{c}=\pi/2$ (black dashed line), the interface length minimizing $E_{\mathcal{C}}$ vanishes (red half-disk).
Simulations comparing ordered and disordered initial cell packings. The total extent of tissue shape change depends on the initial degree of order in the cell packing (cf. Fig. 2). Starting from a hexagonal packing, cell rearrangements are initially well-coordinated leading to coherent tissue flow. As order is lost, tissue flow stalls. The number of cells and the initial tension anisotropy, as well as all other parameters, are the same for both simulations. The simulations were carried out with free boundary conditions. The top row shows the cell arrays, colored by cell coordination number (hexagons=white), bottom row the underlying tension triangulation.
Simulations comparing saturating and winner-takes-all tension feedback. Comparison of saturating (left) and winner-takes-all (right) tension feedback (cf. Fig. 5). Both simulations were initialized with identical initial conditions (an ordered cell array with 20% tension anisotropy). All other parameters are the same for both simulations. The top row shows the cell arrays, colored by cell coordination number (hexagons=white), bottom row the underlying tension triangulation. Saturating feedback leads to little large-scale shape change and creates tension cables.
Simulation combining active and passive tissue regions. The embryo is modeled by a rectangular tissue patch with slip walls on the top and bottom edge, mimicking the embryo’s cylindrical geometry (cf. Fig. 6). The active and passive tissue regions, indicated by color and shading of the edges (color/solid=active), correspond to the germ band and the amnioserosa, respectively. Left: Cell tessellation with black dots indicating the cells along the slip wall (restricted to move horizontally along the “AP axis”). Right: underlying tension triangulation. The tension triangulation is initialized with the experimentally observed initial tension anisotropy which breaks rotational symmetry and aligns the tissue extension along the “AP axis”. Colored dots follow a selected subset of cells to highlight rearrangements.
Movie 5: Simulation of convergent extension without cell rigidity and without DV modulation of activity. Simulation of tissue scale model of convergent extension in an embryo-like geometry for a control and two “mutant” configurations, corresponding Fig. 7. Left: “Control”, as in Movie 3. Middle: Active tissue with vanishing shear modulus $\mu_{a}=0$. Right: No DV modulation, i.e. no soft, passive region. All other parameters are identical. The top shows the cell array, bottom underlying tension triangulation. Active and passive tissue regions are indicated by color and shading of the edges (color/solid=active). In all three conditions, active T1s drive elongation of the tension triangulation, while the convergent extension of the cell array only occurs for the “control” condition.
Figure S1: Coordination number and hexatic order of cells. A Initially (0 min), the majority of cells are hexagonal (light gray). During GBE (25–50 min), topological defects proliferate, resulting in increasing numbers of non-hexagonal cells. Semitransparent gray overlay marks cells that invaginate. B and B’ Fraction of hexagonal cells as a function of time (B) and histogram of cell-coordination number (B’) in the lateral ectoderm. The histograms show pooled data from the time periods 0–15 min and 45–50 min highlighted in B. Lines indicate the coordination number distributions for Voronoi tessellations generated from three different random point processes: hard disk (packing fraction 0.54 near the theoretical maximum for sequential placement [Feder1980]), Ginibre [Ginibre1965], and Poisson. C Hexatic order is low, and shows no long-range correlation. It further decreases as GBE progresses. Hue and brightness code for the phase ($\mathrm{arg}\,\psi_{6}$) and magnitude ($|\psi_{6}|$) of the hexatic order parameter, respectively.
Figure S2: Comparison of local vs regional tension inference. A Local tension inference at a single vertex through the complementarity of angles at a single vertex (left) and in the corresponding tension triangle (right). The relative tensions (edge lengths of the tension triangle) are found via the law of sines. B Regional tension inference for a patch of three cells. The interfaces (and corresponding tensions) shared between them are highlighted by bold lines. The triangulation property of the tensions yields one additional constraint for each cell because for each cell there is a loop in the tension triangulation which has to close. One such loop is highlighted. These additional constraints make the regional inference more robust. C Comparison of the LTC distributions for regional and local tension inference in the Drosophila germ band. Distributions are from data collected over the indicated time intervals. Germ band elongation starts at ca. 25 min. The last row shows single-vertex inference based on vertex positions smoothed using a moving average (1.25 min interval). D Mean anisotropy magnitude (top) and LTC phase (bottom) for the three inference variants. Note the systematically lower LTC phase, i.e. lower fraction of tension bridges, found using single-vertex inference.
Figure S3: LTC distribution for the hard-disk Delaunay family of random triangulation. A Median LTC parameter magnitude and phase vs hard disk packing fraction, $\rho$. B-C LTC histograms for two different packing fractions show significant differences in the LTC distribution. Note that topological statistics such as coordination numbers are much less sensitive to $\rho$. LTC distributions were computed from cell arrays with $4000$ cells.
Figure S4: Illustration of LTC parameter space in terms of internal triangle angles $(\alpha,\beta,\gamma)$. These angles lie in a simplex defined by $\alpha,\beta,\gamma\geq 0$ and $\alpha+\beta+\gamma=\pi$. The simplex has a 6-fold symmetry due to permutations of the angles; its fundamental domain is mapped into the LTC space spanned by $\tilde{s},\tilde{\psi}$ defined in (7). The Delaunay threshold $\max\{\alpha,\beta,\gamma\}\leq\pi$ is indicated as a white dashed line.
Figure S5: Additional quantifications of cell-packing order and tissue deformation. A As cell-packing order is lost due to rearrangements, the cell-shape index $s$ converges to that of a random Voronoi tessellation $s_{\mathrm{V}}$ (dashed black line). Note that the critical shape index for the rigidity transition in the “classical” vertex model [Bi.etal2015], $s_{0}^{*}\approx 3.81$ (dot-dashed orange line), does not feature in our results. Lines show the median, shaded bands indicate quartiles. B Hexagon fraction, $p_{6}$, and tension anisotropy $||\langle\sqrt{2}\tilde{Q}\rangle||$ decrease as cells rearrange and eventually converge to a fixed point where anisotropy vanishes and the hexagon fraction matches a random Voronoi tessellation ($\sim 0.4$). At this point, convergent-extension flow, as measured by the remaining change in tissue aspect ratio $a_{\mathrm{final}}/a(t)$, stalls (cf. Figs. 2C and 2E). The trajectories approximately collapse onto a plane, suggesting that a linear relationship can predict the convergent extension in terms of $p_{6}$ and $||\langle\sqrt{2}\tilde{Q}\rangle||$; see (S2). C Correlation between predicted and observed change in aspect ratio. Data points are from simulations with initial conditions, spanning the entire parameter diagram in (A), showing that the prediction works across many different initial conditions. Points with the same color indicate different time points from the same simulation.
Figure S6: Tension and length dynamics of intercalating interface for T1 events in winner-takes-all and saturating feedback simulations. The shaded area indicates standard deviations. Time $t=0\mathrm{min}$ indicates the moment of the T1 transition. Winner-takes-all feedback matches the corresponding experimental data well (see the companion paper [Brauns.etal2024elife]), while saturating feedback does not, featuring a rapid instead of gradual collapse and reemergence of interfaces. Interface length is measured relative to the average interface length $\overline{\ell}$. T1s occur at a non-zero length due to a numerical cutoff.
Figure S7: Final tissue shapes (40 min simulation time) for the alternative tissue dynamics models discussed in Sec. 2D. Simulations were initialized with an “ordered” tension triangulation and initial anisotropy $s=0.2$. Cells are colored by coordination number (same color code as in Fig. 2A).
Figure S9: Dependence of model behavior on parameters. Each row of plots shows the dependence on one parameter or modeling choice. The three columns follow the same convention as Fig. S8B–D. The passive relaxation time $\tau_{p}$, the balancing time $\tau_{\mathrm{balance}}$, the feedback cutoff, and the initial condition vertex noise are explained in more detail in SI Sec. 5.
Figure S10: Simulations with blocked tissue extension. A Snapshots from a simulation where tissue extension is blocked by slip walls along the AP boundaries of the tissue patch. This mimics experiments where the germ band is blocked from extending via cauterization of the ventral tissue close to the posterior pole, or mutations in the posterior midgut [Collinet.etal2015]. B In simulations where net tissue deformation is blocked (“Toll[RM9]”, “Cauterized”), the interface length dynamics during active T1s is qualitatively similar to simulations of a freely deforming tissue (“WT”). C Orientation of collapsing and emerging interfaces showing that only the orientation of the latter is qualitatively affected by tissue-scale modulation of activity. The orientation of emerging interfaces is slightly biased along the DV direction in the absence of a passive tissue (“Toll[RM9]”, see Fig. 7B for snapshots), while it is biased in AP direction in the “WT” case (see Fig. 6C for snapshots). D When tissue extension is blocked (“Toll[RM9]”) cell rearrangements are compensated by cell elongation leading to a build-up of elastic energy.
Figure S12: Dependence of convergent extension on initial order and anisotropy for vanishing shear rigidity. Heatmap of tissue elongation (square root of aspect ratio change) for vanishing cell shear rigidity, $\mu=0$. (Compare to Fig. 2D where $\mu=1$ was used.) For lower shear rigidity, less convergent extension takes place, as shown in Fig. 7F. The overall structure of the phase diagram remains similar, with weaker dependence on initial order. Note that the simulation was performed with free boundary conditions, i.e. there is no external resistance to tissue deformation. An adjacent passive tissue with non-zero shear modulus would suppress convergent extension further (cf. Fig. 7E).
Figure S13: Cell-level shear modulus determines tissue-level shear modulus. A Large-scale tissue shears are the lowest energy isogonal modes. Reference configuration (minimum energy for a given tension triangulation, left) and deformed configuration (right, outline shows reference configuration), deformed according to the isogonal Hessian eigenvector of one of the two lowest eigenvalues in an example tissue patch. B Tissue level “isogonal shear modulus” (see (S21) for a definition) is proportional to cell level shear modulus. Tissue level isogonal shear modulus vs cell-level shear modulus for the tissue patch shown in A.
Figure S14: Absence of energy barrier to intercalations A Length of the central interface in a quartet for a prescribed isosceles tension triangle parametrized by the angle $\beta$ (the isosceles property implies $\alpha=(\pi-\beta)/2$). B The solid line shows the interface length in the Voronoi tessellation $\ell_{\mathrm{ref}}$ (cf. (8)). The interface length obtained by minimizing elastic energy (3) with isotropic target shape (dot-dashed line) closely follows the Voronoi length and vanishes at the same critical angle. By contrast, minimizing the “area–perimeter” energy of the vertex-model $E_{\mathrm{cell}}\sim(A-A_{0})^{2}+(P-P_{0})^{2}$ in the solid phase, $P_{0}/\sqrt{A_{0}}<3.81$, gives an interface length (dashed black line) that vanishes only for $\beta\to\pi$ and the energy diverges in this limit (dashed green line).
Time stepping Choose time-step size $dt$. For $i=1,...,N_{\mathrm{steps}}$ do: 1. Update intrinsic tensions (positive feedback Euler step): $\tilde{T}_{ij}\mapsto\tilde{T}_{ij}+f(\tilde{T}_{ij},...)dt$. $f$ is the feedback RHS. 2. Update tension triangulation vertices (least-squares fit to intrinsic tensions): $\mathbf{t}_{i}\mapsto\mathrm{argmin}\frac{1}{2N_{\mathrm{edges}}}\sum_{ij}(|\mathbf{t}_{i}-\mathbf{t}_{j}|-\tilde{T}_{ij})^{2}+\frac{1}{2N_{\mathrm{tri}}}\ \sum_{ijk}\mathrm{Pen}(A_{ijk},A_{0})$ 3. Update intrinsic tensions (flatness feedback Euler step): $\tilde{T}_{ij}\mapsto\tilde{T}_{ij}-\tau_{\mathrm{balance}}^{-1}(|\mathbf{t}_{i}-\mathbf{t}_{j}|-\tilde{T}_{ij})dt$ 4. Update cell vertices by angle-constrained energy minimization: $\mathbf{r}_{ijk}\mapsto\mathrm{argmin}\frac{1}{2N_{\mathrm{tri}}}\sum_{ij}T_{ij}\overline{\ell}(1-\hat{n}_{ij}\cdot\mathbf{r}_{ij})+\frac{\varepsilon}{2N_{\mathrm{cells}}}\sum_{i}E_{\mathcal{C},i}(S_{i})$ 5. Update topology: flip all edges of length $\ell_{ij}<\ell_{\mathrm{min}}$, update passive tensions $T_{p,ij}$ and intrinsic tensions $\tilde{T}_{ij}$ in flipped edges. 6. Log: save topology, tension vertices, cell vertices, and T1-transitions.
Figure S15: Simulation workflow showing a summary of the simulation code.
@@ 6 Movie Captions
Passive region shear moduli $\mu_{p},\>\lambda_{p}$ $0.2$ Higher shear resistance of passive tissue lowers the net amount of CE.
@@
Table 1: Parameters of tissue scale simulation. A value in the second-to-last column indicates a “default” value which was used in all simulations except otherwise indicated. $\overline{T}$ is the overall tension scale (average edge tension at the initial condition), and $\ell_{0}$ is the overall length scale (average edge length at the initial condition). Net CE denotes the amount of convergent extension as tissue flow saturates.